Q4. Convert the following linear programming problem into dual


problem.


Maximise


Z = 22x1 + 25x2 +19x3


Subject to:


18x1 + 26x2 + 22x3 ≤ 350


14x1 + 18x2 + 20x3 ≥180


17x1 + 19x2 + 18x3 = 205


x1, x2, x3 ≥ 0



1
Expert's answer
2022-03-08T21:20:01-0500

The primal linear programming problem is


"\text{Maximise~} Z = 22 x_{1} + 25 x_{2} + 19 x_{3}\\
\text{subject to}\\
\begin{aligned}
&18 x_{1}+26 x_{2}+22 x_{3} \leq 350 \\
&14 x_{1}+18 x_{2}+20 x_{3} \geq 180 \\
&17 x_{1}+19 x_{2}+18 x_{3}=205 \\
&\text { and } x_{1}, x_{2}, x_{3} \geq 0
\end{aligned}"


Since the second constraint is of "``\ge"" type, we convert it into "``\le"" by multiplying it by -1.


"\text{Maximise~} Z = 22 x_{1}+25 x_{2}+19 x_{3}\\
\text { subject to } \\
\begin{aligned}
18 x_{1}+26 x_{2}+22 x_{3} &\leq 350\\
-14 x_{1}-18 x_{2}-20 x_{3} &\leq-180 \\
17 x_{1}+19 x_{2}+18 x_{3} &=205\\
\end{aligned}\\
\text{and~} x_{1}, x_{2}, x_{3} \geq 0"


The dual of the given linear programming problem is


"\text{Minimise } Z^*= 350 y_{1}-180 y_{2}+205 y_{3}\\
\text{subject to}\\
\begin{aligned}
18 y_{1}-14 y_{2}+17 y_{3} &\geq 22 \\
26 y_{1}-18 y_{2}+19 y_{3} &\geq 25 \\
22 y_{1}-20 y_{2}+18 y_{3} &\geq 19 \\
\text { and } y_{1}, y_{2} &\geq 0, y_{3} \text { unrestricted in sign }
\end{aligned}"


Since the third constraint in the primal is equality, the corresponding dual variable "y_{3}" will be unrestricted in sign.